# Initialization using Update Approximation is a Silver Bullet for Extremely Efficient Low-Rank Fine-Tuning

### v3

@@ Abstract
Low-rank adapters have become standard for efficiently fine-tuning large language models (LLMs), but they often fall short of achieving the performance of full fine-tuning. We propose a method, LoRA Silver Bullet or LoRA-SB, that approximates full fine-tuning within low-rank subspaces using a carefully designed initialization strategy. We theoretically demonstrate that the architecture of LoRA-XS—which inserts a learnable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed—provides the precise conditions needed for this approximation. We leverage its constrained update space to achieve optimal scaling for high-rank gradient updates while removing the need for hyperparameter tuning. We prove that our initialization offers an optimal low-rank approximation of the initial gradient and preserves update directions throughout training. Extensive experiments across mathematical reasoning, commonsense reasoning, and language understanding tasks demonstrate that our approach exceeds the performance of standard LoRA while using 27-90 times fewer learnable parameters, and comprehensively outperforms LoRA-XS. Our findings establish that it is possible to simulate full fine-tuning in low-rank subspaces, and achieve significant efficiency gains without sacrificing performance. Our code is publicly available at https://github.com/RaghavSinghal10/lora-sb.
Machine Learning, ICML
@@ 1 Introduction
Pre-trained language models have become central to natural language processing, achieving state-of-the-art performance across diverse tasks (Radford et al., 2021; Kirillov et al., 2023; Achiam et al., 2023). While these models excel at general-purpose capabilities (Bubeck et al., 2023; Hao et al., 2022), adapting them to specific downstream tasks often requires fine-tuning (FT). Although in-context learning (Brown et al., 2020; Radford et al., 2019) has gained popularity for its simplicity, it falls short in both performance and efficiency compared to FT (Liu et al., 2022). At the same time, full FT, while highly effective, is computationally expensive and impractical at scale.
@@
Parameter-efficient fine-tuning (PEFT) has become vital for adapting large language models (LLMs) under computational constraints. Low-rank methods like LoRA (Hu et al., 2021) address this by reducing learnable parameters via low-rank updates, sparking advancements in optimization, initialization, structured matrices, and adaptive rank selection (Zhang et al., 2023; Wang et al., 2024b, a). However, these methods face trade-offs: either retain many parameters to match full FT or sacrifice performance for extreme efficiency (Hu et al., 2021; Ding et al., 2023; Wang et al., 2024b). This raises a critical question: Can we design low-rank methods that achieve full FT-level performance while drastically reducing parameter counts?
@@
Figure 1: LoRA-SB. LoRA-XS (Bałazy et al., 2024) reduces parameters compared to LoRA (Hu et al., 2021) by inserting a learnable $r\times r$ matrix $R$ between $B$ and $A$, while keeping other matrices fixed, leading to $W=W_{0}+sBRA$. Our method, LoRA-SB, uses the same architecture. We find that updating $R$ using its gradients $g^{R}$ is equivalent to updating the full FT matrix $W$ with an equivalent gradient $\tilde{g}_{SB}=sBg^{R}A$. We initialize $B$, $R$, and $A$ such that the equivalent gradient $\tilde{g}_{SB}$ provably best approximates the full FT gradient $g$ in low rank subspaces at each step. In essence, we simulate the entire full FT process optimally within low-rank subspaces by utilizing only the first gradient $g_{1}$ (shown in green) from full FT.
@@
Low-rank decomposition methods operate on a fundamental premise: FT requires learning only a low-rank update to the pre-trained weights. Some theoretical work extends this hypothesis, suggesting that methods like LoRA can learn any low-rank approximation of the full FT gradient. However, the gradients computed by these methods do not inherently possess this property. For instance, LoRA’s gradients need explicit optimization at each step to better approximate the full FT gradient (Wang et al., 2024b). Additionally, initialization has emerged as a critical factor in low-rank adaptation, as highlighted by recent works like Pissa-LoRA (Meng et al., 2024) and LoRA-GA (Wang et al., 2024a).
@@
We analyze these limitations in the context of the architecture of LoRA-XS (Bałazy et al., 2024)—which inserts a learnable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed—and demonstrate that these challenges are even more pronounced. While exploring solutions inspired by LoRA-based methods, we discover a remarkable property unique to LoRA-XS: through careful initialization of $A$ and $B$, we can simulate the full FT optimization in low rank subspaces through entire training, as shown in Figure 1. Our initialization provides optimal scaling for approximating high-rank full FT gradients and eliminates need for scaling the hyperparameter $\alpha$.
Our key contributions are:
@@
  • We theoretically formalize the limitations of LoRA-XS, showing how its constrained update space leads to suboptimal gradient approximation, initialization sensitivity, and hyperparameter dependence.
@@
  • We propose an initialization strategy derived from approximating the first step of full FT, proving it provides optimal low-rank approximation of the initial gradient and preserves update directions throughout training.
@@
  • We prove that our initialization makes gradient optimization hyperparameter-independent and guarantees convergence by maintaining orthonormal bases, eliminating the need for any tuning of the scaling factor.
@@
  • Through extensive experiments on $4$ models across $16$ datasets covering mathematical reasoning, commonsense reasoning, and language understanding tasks, we demonstrate that our method surpasses LoRA’s performance while using 27-90x less learnable parameters, and comprehensively outperforms LoRA-XS.
2 Related Work
Parameter-Efficient Fine-Tuning (PEFT). PEFT methods have become essential for adapting large pre-trained models under computational constraints. Early techniques like AdapterFusion (Pfeiffer et al., 2021) and Prefix-Tuning (Li & Liang, 2021) enabled task-specific adaptation with minimal parameter updates. Advances like soft prompts (Lester et al., 2021) further reduced trainable parameter counts while maintaining strong performance. Recent approaches have explored operating directly on model representations (Wu et al., 2024).
Low-Rank Decomposition Methods. LoRA (Hu et al., 2021) demonstrated that weight updates during FT could be efficiently approximated using low-rank matrices, drastically reducing parameter counts. Building on this insight, variants such as QLoRA (Dettmers et al., 2023) and AdaLoRA (Zhang et al., 2023) extended the paradigm through quantization and adaptive allocation strategies. The applicability of low-rank techniques has also been explored in pretraining with GaLore (Zhao et al., 2024) and ReLoRA (Lialin et al., 2023), highlighting the versatility of low-rank adaptation methods. LoRA-based methods have also been applied in other domains, such as efficient federated FT (Sun et al., 2024; Singhal et al., 2025).
Enhancing LoRA Performance. Recent efforts have focused on optimizing LoRA’s performance. PiSSA (Meng et al., 2024) demonstrated improvements by initializing matrices with principal components of pre-trained weights. LoRA-Pro (Wang et al., 2024b) and LoRA-GA (Wang et al., 2024a) improved gradient approximation, aligning low-rank updates more closely with full FT. Methods like DoRA (Liu et al., 2024) and rsLoRA (Kalajdzievski, 2023) introduced decomposition-based and scaling stabilization techniques to enhance learning stability and expand LoRA’s utility.
Improving Efficiency in LoRA Variants. Efficiency-focused innovations have pushed LoRA toward more parameter savings. LoRA-XS (Bałazy et al., 2024) achieves this by inserting a small trainable weight matrix into frozen low-rank matrices. VeRA (Kopiczko et al., 2024) shares low-rank matrices across layers, relying on scaling vectors for task-specific adaptation. Tied-LoRA (Renduchintala et al., 2024) leverages weight tying to reduce parameter usage at higher ranks, while HydraLoRA (Tian et al., 2024) introduces an asymmetric architecture for improvement.
@@ 3.1 Preliminaries
where $B$ and $A$ are fixed, and only $R\in\mathbb{R}^{r\times r}$ is trainable, reducing the number of parameters to $r^{2}$. We denote the full FT gradient as $g=\frac{\partial L}{\partial W}$, and the LoRA-XS gradient as $g^{R}_{\text{LoRA-XS}}=\frac{\partial L}{\partial R}$, where $L$ is the loss function.
@@ 3.2 Motivation
LoRA-XS (Bałazy et al., 2024) has significantly fewer learnable parameters than LoRA but performs suboptimally. LoRA-XS’s architecture causes constraints on the type of updates it can learn. The subspace of learned updates is characterized in Lemma 1. This implies that while $\Delta W$ is constrained to be rank $\leq r$, it also needs to have column and row spaces defined by those of $B$ and $A$, respectively. In contrast, LoRA can learn any update $\Delta W$ as long as $\text{rank}(\Delta W)\leq r$. Thus, the low expressivity of LoRA-XS as compared to LoRA can account for the performance drop.
@@
Let $\Delta W$ be an update learned with LoRA-XS. Then, the set of all possible $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, is given as: $\{M\in\mathbb{R}^{m\times n}|\text{Col}(M)\subseteq\text{Col}(B)$ $\wedge\text{Row}(M)\subseteq\text{Row}(A)\},$ where $\text{Col}(M)$ and $\text{Row}(M)$ are column and row spaces of matrix $M$ respectively.
@@
1) Inadequate Gradient Approximation: LoRA optimization is mathematically equivalent to full FT using a constrained low-rank gradient. The gradient of LoRA does not optimally approximate the full gradient, and needs to be tuned at each step. LoRA-Pro (Wang et al., 2024b) finds that this results in suboptimal performances, and provides a closed form solution to optimize the gradients. In LoRA-XS, the gradient updates are restricted to an even more constrained low-rank space since $A$ and $B$ are fixed. We posit that the limitation becomes particularly severe when the ideal updates lie outside the space spanned by fixed $A$ and $B$, and consequently has a larger impact on performance.
2) Suboptimal Initialization: While initialization impacts all low-rank methods, it becomes critical in LoRA-XS where $A$ and $B$ are frozen. Unlike LoRA where poor initialization can be compensated through training, LoRA-XS relies entirely on its initial subspace defined by $A$ and $B$. Consider the zero initialization of the $B$ matrix, for example. While LoRA may experience some performance degradation in this case (Wang et al., 2024a; Meng et al., 2024), the ideal low-rank update $\Delta W$ can still be reached through gradient descent. In fact, zero initialization for the $B$ matrix is commonly used, including in the original LoRA paper (Hu et al., 2021). However, in LoRA-XS, this results in no learning, as the product $BRA$ remains zero. LoRA-XS uses the most significant subspaces spanned by the columns of pre-trained weights for initialization, inspired by Meng et al. (2024). This initialization is not aligned well with FT because it fails to capture the specific subspaces relevant to the FT task.
3) Hyperparameter Sensitivity: The scaling factor $s$, present in almost every LoRA based FT method requires tuning to maintain stability during training. This factor acts as a bridge between the low-rank and full-rank spaces, compensating for the dimensional mismatch in gradients. Poor tuning of $s$ can lead to unstable training or slow convergence, e.g., see rsLoRA (Kalajdzievski, 2023), adding complexity and potentially limiting practical deployment.
@@ 3.3 Approximation of the full FT gradient
As mentioned, LoRA optimization is mathematically equivalent to full FT using a constrained low-rank gradient. However, the update generated using the gradients of LoRA does not result in the same update which the low-rank gradient would have generated. The following holds true for LoRA-XS as well. To understand this, let us look at the change in weight $W$ and its relationship with changing of low-rank matrix $R$, which can be simply given by $\mathrm{d}W=-sB(\mathrm{d}R)A.$ This implies that updating $R$ with gradient $g^{R}$ is equivalent to updating $W$ with low rank equivalent gradient $\tilde{g}$ in full FT as described in Definition 1.
@@
We define the equivalent gradient as: $\tilde{g}=sBg^{R}A$ where $g^{R}$ is the gradient of $L$ with respect to $R$.
@@
The gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{\top}gA^{\top}$
@@
See Appendix A.2. ∎
We can now formulate our objective to minimize the distance between the equivalent gradient and the full gradient. We do not have access to the full FT gradient $g$ during LoRA-XS based FT. Thus we need to find the ideal gradient with respect to $R$, given by $g^{R}$, and subsequently the optimal approximation $\tilde{g}$, in terms of the gradient which is available to us during training: $g^{R}_{LoRA-XS}$. Fortunately, this optimization problem admits a closed-form solution independent of $g$ as described in Theorem 3.
@@
The optimal solution for the objective $\text{min}_{g^{R}}||\tilde{g}-g||^{2}_{F}$, such that $\tilde{g}=sBg^{R}A$, is: $g^{R}=\dfrac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{LoRA-XS}(AA^{\top})^{-1}$ (4)
@@
See Appendix A.3. ∎
The closed-form solution in Theorem 3 solves the optimization problem $\text{min}_{g^{R}}||\tilde{g}-g||^{2}_{F}$, but by itself doesn’t ensure the loss will decrease when updating $R$. Through Theorem 4, we prove that the change in loss is non-positive ($\Delta L\leq 0$). This property is fundamental to optimization as it guarantees consistent loss minimization throughout training.
@@
Consider the update for matrix $R$ using the solution derived in Theorem 3: $R\leftarrow R-\eta g^{R}$ where $\eta>0$ is the (sufficiently small) learning rate. This update guarantees a reduction in the loss $\Delta L$, given by: $\Delta L=-\eta\langle g^{R}_{LoRA-XS},g^{R}\rangle_{F}+o(\eta)\leq 0.$
@@ 3.4 Initialization using update approximation
This better captures the general direction of adaptation required for the target task while being less sensitive to individual sample variations. We can then use truncated SVD to obtain a low-rank approximation of $\Delta W_{\text{avg}}$, and express it as $sBRA$. There exist infinite combinations of $B$ and $A$ which can obey this relationship. For instance, we can initialize $B$ and $A$ as $US$ and $V^{\top}$ and keep $R$ as $I/s$. This is equivalent to the $B$ and $A$ initialization in LoRA-XS but by approximating the update rather than the pre-trained matrix. We note that the above process can be computed for any optimizer, by approximating the corresponding first step. We compute this specifically for AdamW since we use it.
3.5 Hyperparameter independence
The hyperparameter $\alpha$ is used in LoRA and other decomposition-based method to tackle the issue of instability caused to improper scaling of the updates. The gradient scaling is accounted for, by adding a hyperparameter to normalize the updates. The importance of scaling is shown in methods like rank stabilization (Kalajdzievski, 2023). However, the full FT gradient $g$ needs no such tuning. We claim that approximating the full FT gradient removes the need for introducing a scaling factor, as shown in Theorem 5.

### v4

@@ Abstract
Low-rank adapters have become standard for efficiently fine-tuning large language models, but they often fall short of achieving the performance of full fine-tuning. We propose a method, LoRA Silver Bullet or LoRA-SB, that approximates full fine-tuning within low-rank subspaces using a carefully designed initialization strategy. We theoretically demonstrate that the architecture of LoRA-XS, which inserts a learnable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed, provides the precise conditions needed for this approximation. We leverage its constrained update space to achieve optimal scaling for high-rank gradient updates while removing the need for scaling factor tuning. We prove that our initialization offers an optimal low-rank approximation of the initial gradient and preserves update directions throughout training. Extensive experiments across mathematical reasoning, commonsense reasoning, and language understanding tasks demonstrate that our approach exceeds the performance of LoRA (and baselines) while using 27-90 times fewer learnable parameters, and comprehensively outperforms LoRA-XS. Our findings establish that it is possible to simulate full fine-tuning in low-rank subspaces, and achieve significant parameter efficiency gains without sacrificing performance. Our code is publicly available at: https://github.com/CERT-Lab/lora-sb.
@@ 1 Introduction
Pre-trained language models have become central to natural language processing, achieving state-of-the-art performance across diverse tasks (35, 21, 1). While these models excel at general-purpose capabilities (4, 14), adapting them to specific downstream tasks often requires fine-tuning (FT). At the same time, full FT, while highly effective, is computationally expensive and impractical at scale.
@@
Parameter-efficient fine-tuning (PEFT) has become vital for adapting large language models (LLMs) under computational constraints. Low-rank methods like LoRA (17) address this by reducing learnable parameters via low-rank updates, sparking advancements in optimization, initialization, structured matrices, and adaptive rank selection (52, 46, 45). However, these methods face trade-offs: either retain many parameters to match full FT or sacrifice performance for extreme efficiency (17, 10, 46). This raises a critical question: Can we design low-rank methods that achieve full FT-level performance while drastically reducing parameter counts?
@@
Figure 1: LoRA-SB. LoRA-XS (2) reduces parameters compared to LoRA (17) by inserting a learnable $r\times r$ matrix $R$ between $B$ and $A$, while keeping other matrices fixed, leading to $W=W_{0}+sBRA$. Our method, LoRA-SB, uses the same architecture. We find that updating $R$ using its gradients $g^{R}$ is equivalent to updating the full FT matrix $W$ with an equivalent gradient $\tilde{g}_{SB}=sBg^{R}A$. We initialize $B$, $R$, and $A$ such that the equivalent gradient $\tilde{g}_{SB}$ provably best approximates the full FT gradient $g$ in low rank subspaces at each step. In essence, we simulate the entire full FT process optimally within low-rank subspaces by utilizing only the first full FT gradient $g_{1}$.
@@
Low-rank decomposition methods operate on a fundamental premise: FT requires learning only a low-rank update to the pre-trained weights. However, the gradients computed by these methods do not inherently possess this property. For instance, LoRA’s gradients need explicit optimization at each step to better approximate the full FT gradient (46). Additionally, initialization has emerged as a critical factor in low-rank adaptation, as highlighted by recent works like PiSSA-LoRA (30) and LoRA-GA (45).
@@
We analyze these limitations in the context of the architecture of LoRA-XS (2), which inserts a learnable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed, and demonstrate that these challenges are even more pronounced. While exploring solutions inspired by LoRA-based methods, we discover a remarkable property unique to LoRA-XS: through careful initialization of $A$ and $B$, we can simulate the full FT optimization in low rank subspaces through entire training, as shown in Figure 1. Our initialization provides optimal scaling for approximating high-rank full FT gradients and eliminates need for tuning the hyperparameter $\alpha$. The peak memory usage of LoRA-SB never exceeds that of LoRA or other baselines, and its training-time overhead relative to LoRA is negligible ($\approx 1.1\%-1.3\%$). Our key contributions are:
@@
  • We formalize the limitations of LoRA-XS, showing how its constrained update space leads to suboptimal gradient approximation, initialization sensitivity, and scaling dependence.
@@
  • We propose an initialization strategy derived from using the first step of full FT, which provides an optimal approximation of the initial gradient and preserves update directions throughout.
@@
  • We prove our initialization makes gradient optimization scaling-independent and guarantees convergence by maintaining orthonormal bases, eliminating need for tuning the scaling factor $\alpha$.
@@
  • Through extensive experiments on $4$ models across $16$ datasets covering mathematical reasoning, commonsense reasoning, and language understanding, we demonstrate that LoRA-SB surpasses LoRA while using 27-90x less learnable parameters, and comprehensively outperforms LoRA-XS.
@@ 2.1 Preliminaries
where $B$ and $A$ are fixed, and only $R\in\mathbb{R}^{r\times r}$ is trainable, reducing the number of parameters to $r^{2}$. We denote the full FT gradient: $g=\frac{\partial L}{\partial W}$; LoRA-XS gradient: $g^{R}_{\text{LoRA-XS}}=\frac{\partial L}{\partial R}$; $L$ is the loss function.
@@ 2.2 Motivation
LoRA-XS (2) has significantly fewer learnable parameters than LoRA but performs suboptimally. LoRA-XS’s architecture causes constraints on the type of updates it can learn. The subspace of learned updates is characterized in Lemma 1. This implies that while $\Delta W$ is constrained to be rank $\leq r$, it also needs to have column and row spaces defined by those of $B$ and $A$, respectively. In contrast, LoRA can learn any update $\Delta W$ as long as $\text{rank}(\Delta W)\leq r$. Thus, the low expressivity of LoRA-XS as compared to LoRA can account for the performance drop.
@@
Let $\Delta W$ be an update learned with LoRA-XS. Then, the set of all possible $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, is given as: $\mathcal{W}_{LoRA-XS}=\{M\in\mathbb{R}^{m\times n}|\text{Col}(M)\subseteq\text{Col}(B)\wedge\text{Row}(M)\subseteq\text{Row}(A)\},$ where $\text{Col}(M)$ and $\text{Row}(M)$ are column and row spaces of matrix $M$ respectively.
@@
1) Inadequate Gradient Approximation: LoRA optimization is mathematically equivalent to full FT using a constrained low-rank gradient. The gradient of LoRA does not optimally approximate the full gradient, and needs to be tuned at each step. LoRA-Pro (46) finds that this results in suboptimal performances, and provides a closed form solution to optimize the gradients. In LoRA-XS, the gradient updates are restricted to an even more constrained low-rank space since $A$ and $B$ are fixed. We posit that the limitation becomes particularly severe when the ideal updates lie outside the space spanned by fixed $A$ and $B$, and consequently has a larger impact on performance.
2) Suboptimal Initialization: While initialization impacts all low-rank methods, it becomes critical in LoRA-XS where $A$ and $B$ are frozen. Unlike LoRA where poor initialization can be compensated through training, LoRA-XS relies entirely on its initial subspace defined by $A$ and $B$. Consider the zero initialization of the $B$ matrix, for example. While LoRA may experience some performance degradation in this case (45, 30), the ideal low-rank update $\Delta W$ can still be reached through gradient descent. In fact, zero initialization for the $B$ matrix is commonly used, including in the original LoRA paper (17). However, in LoRA-XS, this results in no learning, as the product $BRA$ remains zero. LoRA-XS uses the most significant subspaces spanned by the columns of pre-trained weights for initialization, inspired by PiSSA (30). This initialization is not aligned well with FT because it fails to capture the specific subspaces relevant to the FT task.
3) Scaling Factor Sensitivity: The scaling factor $s$, present in almost every LoRA based method, requires tuning to maintain stability during training. This factor acts as a bridge between the low-rank and full-rank spaces, compensating for the dimensional mismatch in gradients. Poor tuning of $s$ can lead to unstable training or slow convergence (rsLoRA (20)), adding complexity and potentially limiting practical deployment.
@@ 2.3 Approximation of the full FT gradient
As mentioned, LoRA optimization is equivalent to full FT using a constrained low-rank gradient. However, the update generated using the gradients of LoRA does not result in the same update which the low-rank gradient would have generated. The following holds true for LoRA-XS as well. To understand this, let us look at the change in weight $W$ and its relationship with changing of low-rank matrix $R$, which can be simply given by $\mathrm{d}W=-sB(\mathrm{d}R)A.$ This implies that updating $R$ with gradient $g^{R}$ is equivalent to updating $W$ with low rank equivalent gradient $\tilde{g}$ in full FT (Definition 1).
@@
We define the equivalent gradient in LoRA-XS as: $\tilde{g}=sBg^{R}A$, where $g^{R}$ is the gradient of $L$ with respect to $R$.
@@
The gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{\top}gA^{\top}$.
@@
See Appendix B.2. ∎
We now formulate our objective to minimize the distance between the equivalent gradient and the full gradient. We do not have access to the full FT gradient $g$ during LoRA-XS based FT. Thus we need to find the ideal gradient with respect to $R$, given by $g^{R}$, and subsequently the optimal approximation $\tilde{g}$, in terms of the gradient which is available to us during training: $g^{R}_{LoRA-XS}$. Fortunately, this optimization problem admits a closed-form solution independent of $g$ as described in Theorem 3.
@@
For full-rank AA and $B$ matrices, the optimal solution for the objective
[…]
@@
See Appendix B.3. ∎
[…]
@@
Consider the update for matrix $R$ using the solution derived in Theorem 3:
$R\leftarrow R-\eta g^{R}$, where $\eta>0$ is the (sufficiently small) learning rate. This update guarantees a reduction in the loss $\Delta L$, given by: $\Delta L=-\eta\langle g^{R}_{LoRA-XS},g^{R}\rangle_{F}+o(\eta)\leq 0.$
@@ 2.4 Initialization using update approximation
Since AdamW is used as the optimizer for both full FT and LoRA-SB training, we approximate its first update step using the sign of the summed gradients rather than their raw values (see Appendix C for details). This better captures the direction of adaptation required for the target task while being less sensitive to individual sample variations. We then use truncated SVD to obtain a low-rank approximation of $\Delta W_{\text{avg}}$, and express it as $sBRA$. There exist infinite combinations of $B$ and $A$ which can obey this relationship. For instance, we can initialize $B$ and $A$ as $US$ and $V^{\top}$ and keep $R$ as $I/s$. This is equivalent to the $B$ and $A$ initialization in LoRA-XS but by approximating the update rather than the pre-trained matrix. The above process can be computed for any optimizer, by approximating the corresponding first step. We compute this specifically for AdamW since we use it.
2.5 Scaling Factor independence
The hyperparameter $\alpha$ is used in LoRA and other decomposition-based methods to tackle instability caused to improper scaling of the updates. The gradient scaling is accounted for, by adding a hyperparameter to normalize the updates. The importance of scaling is shown in methods like rank stabilization (20). However, the full FT gradient $g$ needs no such tuning. We claim that approximating the full FT gradient removes the need for introducing a scaling factor, as shown in Theorem 5.